# NRGBoost: Energy-Based Generative Boosted Trees

### v1

@@ 7 Discussion
The proposed DEF models are easy to sample from and require no sampling during training. However, we find that, they require deep trees to model the data well which, in turn, also requires a larger number of trees in the ensemble to regularize. In our experiments we found that the performance of DEF models was often capped by the maximum number of leaves we allowed ($2^{14}$).
In contrast, NRGBoost was able to model the data better while using shallower trees and in fewer number (the models selected on all datasets except for MNIST and CT had $\leq 2^{10}$ leaves). Its main downside is that it can only be sampled from approximately using more expensive MCMC and also requires sampling during the training process. While our fast Gibbs sampling implementation and our proposed sampling approach were able to mitigate the slow training, making these models practical, they can still be cumbersome to use for sampling, particularly if independent samples are required, due to autocorrelation between samples from the same Markov Chain. We argue however that unlike in image or text generation where fast sampling is necessary for an interactive user experience, this can be less of a concern for the task of generating synthetic datasets where the one time cost of sampling is not as important as faithfully capturing the data generating distribution.
A limitation of tree-based models when compared to DL approaches is that they require keeping the full training data in memory. While discretization allows this data to be stored in a compact format, this could still limit the applicability of these methods to very large datasets.
@@ 8 Conclusion
We extend the two most popular tree-based discriminative methods for use in generative modeling. We find that our boosting approach, in particular, offers generally good discriminative performance and competitive sampling performance to more specialized alternatives. We hope that these results encourage further research into generative boosting approaches for tabular data, in particular exploring other applications besides sampling that are enabled by (unnormalized) density models.
@@ Appendix A Theory
Analogous to a Hessian, we can differentiate Equation 17 again along a second independent variation $h$ of $f$ yielding a symmetric bilinear functional which we will write as $\delta^{2}L[f;g,h]$. Note that the first term in equation 16 is linear in $f$ and thus has no curvature, so we only have to consider the log partition function itself:
@@
Note that this functional is negative semi-definite for all $f$, i.e. $\delta^{2}L[f;h,h]\leq 0$, meaning that the log-likelihood is a concave functional of $f$.
@@
As an aside, defining the functional derivative, $\frac{\delta J[f]}{\delta f(\mathbf{x})}$, of a functional $J$ implicitly by:
@@
we can formally define, by analogy with the parametric case, the Fisher Information "Matrix" (FIM) at $f$ as the following bilinear functional of two independent variations $g$ and $h$:
@@
$=\delta^{2}\log Z[f;g,h]\,.$ (34)
The only difference to the second-order variation of 16 computed in equation 22 would be that the expectation is taken under the model distribution, $q_{f}$, instead of the data distribution $p$. However, because the only term in $\log q_{f}(\mathbf{x})$ that is non-linear in $f$ is the log-partition functional, which is not a function of $\mathbf{x}$, this expectation plays no role in the computation and we get the result that the Fisher Information is the same as the negative Hessian of the log-likelihood for these models.
@@ A.1 Application to Piecewise Constant Functions
$1=\sum_{j=1}^{J}\mathbf{1}_{X_{j}}(\mathbf{x})\,.$ (42)
@@
and of the result in equation 41 in the final step.
This result can be used to conduct a line search over the step size using training data and to estimate an increase in likelihood at each round of boosting for the purpose of early stopping, using validation data.
A.2 A Simple Lower Bound on the Likelihood Increase
We can take the results from the last subsection to show that each boosting round improves the likelihood of the model under ideal conditions. Substituting the NRGBoost leaf values $w_{j}=\nicefrac{{P(X_{j})}}{{Q_{f}(X_{j})}}-1$ in Equation 45 we get:
@@ A.2 A Simple Lower Bound on the Likelihood Increase
The second term in Equation 48 can be interpreted as the log of a moment generating function for a centered random variable $W$ that takes values in $\left\{\nicefrac{{P\left(X_{j}\right)}}{{Q_{f}(X_{j})}}-1\right\}_{j\in[1..J]}$ (i.e., the set of the leaf values), with corresponding probabilities $\left\{Q_{f}(X_{j})\right\}_{j\in[1..J]}$. With our proposed regularization approach (see Section 3.3), we limit the ratio $\nicefrac{{P\left(X_{j}\right)}}{{Q_{f}(X_{j})}}$ in any given $X_{j}$ to a maximum value $R$ when partitioning the domain and therefore we have that $W\in[-1,R-1]$.
Since $W$ is bounded, by Hoeffding’s Lemma (1963), it is sub-gaussian with variance proxy $R^{2}/4$ and the log of its moment generating function is thus upper bounded by:
@@
As long as in the current round of boosting a partitioning of the input space can be found that yields a non-zero $\chi^{2}(P\|Q_{f})$, the likelihood is guaranteed to increase as long as we choose a small enough step-size: $\alpha\in\left(0,\frac{8}{R^{2}}\chi^{2}(P\|Q_{f})\right)$. In particular, choosing a step size $\alpha=\frac{4}{R^{2}}\chi^{2}(P\|Q_{f})$ produces an increase in log-likelihood of at least $2\left[\nicefrac{{\chi^{2}(P\|Q_{f})}}{{R}}\right]^{2}$.
We note that this result, while insightful, assumes that one uses the exact probabilities under the model distribution in the NRGBoost update. In practice, we would only be able to approximately estimate these with MCMC. Furthermore, it can only guarantee an increase in the training log-likelihood as we also don’t have access to the probabilities under the data generating distribution. We leave further analysis for future work.
Appendix B Density Estimation Trees
Density Estimation Trees (DET) (Ram & Gray, 2011) model the density function as a piecewise constant function,
@@ Appendix B Density Estimation Trees
where $X_{j}$ are given by a partitioning of the input space $\mathcal{X}$ induced by a binary tree and the $v_{j}$ are the density values associated with each leaf that, for the time being, we will only require to be such that $q(\mathbf{x})$ sums to one.
Ram & Gray (2011) proposes fitting DET models to directly minimize a generative objective, the Integrated Squared Error (ISE) between the data generating distribution, $p(\mathbf{x})$ and the model:
@@
Noting that $q$ is a function as in Equation 51 and that $\bigcup_{j=1}^{J}X_{j}=\mathcal{X}$, we can rewrite this as
@@
Since the first term in the objective does not depend on the model this optimization problem can be further simplified as
@@
One can therefore grow a tree by greedily choosing to split a parent leaf with support $X_{P}$ into two leaves with supports $X_{L}$ and $X_{R}$ so as to maximize the following criterion:
@@
To maximize the likelihood,
@@
rather than the ISE one can use the same approach. Here the optimization problem to solve is:
@@
This is, again, easy to solve for $v_{j}$ since it is separable over $j$ after removing the constraint using Lagrange multipliers. The optimal leaf values and objective are in this case:
@@
The only change is therefore to the splitting criterion which should become:
@@ Appendix C Greedy Tree-Based Multiplicative Boosting
In multiplicative generative boosting an unnormalized current density model, $\tilde{q}_{t-1}(\mathbf{x})$, is updated at each boosting round by multiplication with a new factor $\delta q_{t}^{\alpha_{t}}(\mathbf{x})$:
@@
For our proposed NRGBoost, this factor is chosen in order to maximize a local quadratic approximation of the log likelihood around $q_{t-1}$ as a functional of the log density (see Section 3). The motivation behind the greedy approach of Tu (2007) or Grover & Ermon (2018) is to instead make the update factor $\delta q_{t}(\mathbf{x})$ proportional to the likelihood ratio $r_{t}(\mathbf{x})\coloneqq\nicefrac{{p(\mathbf{x})}}{{q_{t-1}(\mathbf{x})}}$ directly, which under ideal conditions would mean that the method converges immediately when choosing a step size $\alpha_{t}=1$. In a more realistic setting, however, this method has been shown to converge under conditions on the performance of the individual $\delta q_{t}$ as discriminators between real and generated data (Tu, 2007; Grover & Ermon, 2018; Cranko & Nock, 2019).
While in principle this desired $r_{t}(\mathbf{x})$ could be derived from any binary classifier that is trained to predict a probability of a datapoint being generated (e.g., by training it to minimize a strictly proper loss) and Tu (2007) does not prescribe any particular choice, Grover & Ermon (2018) propose relying on the following variational bound of an $f$-divergence to derive an estimator for this ratio:
$D_{f}(P\|Q_{t-1})\geq\sup_{u\in\mathcal{U}_{t}}\left[\mathbb{E}_{\mathbf{x}\sim p}\;u(\mathbf{x})-\mathbb{E}_{\mathbf{x}\sim q_{t-1}}f^{*}(u(\mathbf{x}))\right]\,.$ (62)
Here $f^{*}$ denotes the convex conjugate of $f$. This bound is tight, with the optimum being achieved for $u_{t}^{*}(\mathbf{x})=f^{\prime}(\nicefrac{{p(\mathbf{x})}}{{q_{t-1}(\mathbf{x})}})$, if $\mathcal{U}_{t}$ is capable of representing this function. $\left(f^{\prime}\right)^{-1}(u^{*}_{t}(\mathbf{x}))$ can thus be interpreted as an approximation of the desired $r_{t}(\mathbf{x})$.
Adapting this method to use trees as weak learners can be accomplished by considering $\mathcal{U}_{t}$ in Equation 62 to be defined by tree functions $u=\nicefrac{{1}}{{J}}\sum_{j=1}^{J}w_{j}\mathbf{1}_{X_{j}}$ with leaf values $w_{j}$ and leaf supports $X_{j}$. At each boosting iteration a new tree, $u^{*}_{t}$ can thus be grown to greedily optimize the lower bound in the r.h.s. of Equation 62 and setting $\delta q_{t}(\mathbf{x})=\left(f^{\prime}\right)^{-1}(u^{*}_{t}(\mathbf{x}))$ which is thus also a tree with the same leaf supports and leaf values given by $v_{j}\coloneqq\left(f^{\prime}\right)^{-1}(w_{j})$. This leads to the seaprable optimization problem:
@@
Note that we drop the iteration indices from this point onwards for brevity. Maximizing over $w_{j}$ with the $X_{j}$ fixed we have that $w_{j}^{*}=f^{\prime}\left(\nicefrac{{P(X_{j})}}{{Q(X_{j})}}\right)$ which yields the optimal value
$J^{*}(X_{1},\ldots,X_{j})=\sum_{j}\left[P(X_{j})f^{\prime}\left(\frac{P(X_{j})}{Q(X_{j})}\right)-Q(X_{j})(f^{*}\circ f^{\prime})\left(\frac{P(X_{j})}{Q(X_{j})}\right)\right]$ (64)
In turn, this determines the splitting criterion as a function of the choice of $f$. Finally, the optimal density values for the leaves are given by
@@
It is interesting to note two particular choices of $f$-divergences. For the KL divergence, $f(t)=t\log t$ and $f^{\prime}(t)=1+\log t=\left(f^{*}\right)^{-1}(t)$. This leads to
@@
as the splitting criterion. The Pearson $\chi^{2}$ divergence, with $f(t)=(t-1)^{2}$, leads to the same splitting criterion as NRGBoost. Note however that for NRGBoost the leaf values for the multiplicative update of the density are proportional to $\exp\left(\nicefrac{{P(X_{j})}}{{Q(X_{j})}}\right)$ instead of the ratio directly. Table 6 summarizes these results.
@@
Another interesting observation is that a DET model can be interpreted as a single round of greedy multiplicative boosting starting from a uniform initial model. The choice of the ISE as the criterion to optimize the DET corresponds to the choice of Pearson’s $\chi^{2}$ divergence and likelihood to the choice of KL divergence.

### v2

@@ 6 Discussion
While additive tree ensemble models like DEF require no sampling to train and are easy to sample from, we find that, in practice, they require very deep trees to model the data well and still offer subpar performance for sampling and density estimation.
In contrast, NRGBoost was able to model the data better, while using fewer and shallower trees. Its main drawback is that it can only be sampled from approximately using more expensive MCMC and also requires sampling during the training process. Our fast Gibbs sampling implementation, coupled with our proposed amortized sampling approach, were able to mitigate the slow training, making these models more practical. Unfortunately, they remain cumbersome to use for sampling, as autocorrelation between samples from the same Markov chain makes them inefficient in scenarios requiring independent samples. We argue, however, that unlike in image or text generation, where fast sampling is necessary for an interactive user experience, this can be less of a concern when generating synthetic datasets, where the one-time cost of sampling can be less important than faithfully capturing the data-generating distribution.
One advantage of the energy-based approach that we did not explore is that it allows for arbitrary conditional sampling, requiring only that one clamps the values of conditioning variables during Gibbs sampling. In contrast, other methods, such as diffusion-based approaches, are less flexible, since conditioning variables typically need to be predetermined at training time.
@@ 7 Conclusion
We extend the two most popular tree-based discriminative methods for use in generative modeling. We find that our boosting approach, in particular, offers generally good discriminative performance and competitive sampling performance to more specialized alternatives. We hope that these results encourage further research into generative boosting approaches for tabular data, particularly in exploring other applications beyond sampling that are enabled by density models.
@@ Appendix A Theory
Analogous to a Hessian, we can differentiate Equation 15 again along a second independent variation $h$ of $f$, yielding a symmetric bilinear functional which we will write as $\delta^{2}L[f;g,h]$. Note that the first term in Equation 14 is linear in $f$ and thus has no curvature, so we only have to consider the log-partition function itself:
@@
Note that this functional is negative semi-definite for all $f$. In other words, we have $\delta^{2}L[f;h,h]\leq 0$ for any variation $h$, meaning that the log-likelihood is a concave functional of $f$.
@@
As an aside, defining the functional derivative, $\frac{\delta J[f]}{\delta f(\mathbf{x})}$, of a functional $J$ implicitly by
@@
we can formally define, by analogy with the parametric case, the Fisher information "matrix" at $f$, as the following bilinear functional of two independent variations $g$ and $h$:
@@
$=\delta^{2}\log Z[f;g,h]=-\delta^{2}L[f;g,h]\,.$ (32)
To summarize, we have that the Fisher information is the same as the negative Hessian of the log-likelihood for these models. This is essentially due to the fact that the only term in $\log q_{f}(\mathbf{x})$ that is non-linear in $f$ is the log-partition, which is not a function of $\mathbf{x}$, and therefore it doesn’t matter whether the expectation is taken under $p$ or $q_{f}$.
@@ A.1 Application to Piecewise Constant Functions
$1=\sum_{j=1}^{J}\mathbf{1}_{X_{j}}(\mathbf{x})\qquad\forall\mathbf{x}\in\mathcal{X}\,.$ (40)
@@
and of the result in Equation 39 in the final step. This result can be used to conduct a line search over the step-size using training data and to estimate an increase in log-likelihood at each round of boosting, for the purpose of early stopping, using validation data.
A.2 A Simple Convergence Result
We can take the results from the last subsection to show that each boosting round improves the likelihood of the model under ideal conditions. Substituting the NRGBoost leaf values, $w_{j}=\nicefrac{{P(X_{j})}}{{Q_{f}(X_{j})}}-1$, in Equation 43 we get:
@@ A.2 A Simple Convergence Result
The second term in Equation 46 can be interpreted as the log of a moment generating function for a centered random variable $W$ that takes values in $\left\{\nicefrac{{P\left(X_{j}\right)}}{{Q_{f}(X_{j})}}-1\right\}_{j\in[1..J]}$ (i.e., the set of the leaf values), with probabilities $\left\{Q_{f}(X_{j})\right\}_{j\in[1..J]}$. With our proposed regularization approach, we limit the ratio $\nicefrac{{P\left(X_{j}\right)}}{{Q_{f}(X_{j})}}$ in our choice of $X_{j}$ to a maximum value $R$ and therefore we have that $W\in[-1,R-1]$. Given that $W$ is bounded, by Hoeffding’s Lemma (1963), it is sub-gaussian with variance proxy $R^{2}/4$ and the log of its moment generating function is thus upper bounded by:
@@
As long as, in the current round of boosting, a partitioning of the input space can be found that yields a non-zero $\chi^{2}(P\|Q_{f})$, the likelihood is guaranteed to increase as long as we choose a small enough step size, $\alpha<\frac{8}{R^{2}}\chi^{2}(P\|Q_{f})$. In particular, choosing a step size $\alpha=\frac{4}{R^{2}}\chi^{2}(P\|Q_{f})$ produces an increase in likelihood of at least $2\left[\nicefrac{{\chi^{2}(P\|Q_{f})}}{{R}}\right]^{2}$.
We note that this result, while insightful, assumes that one uses the exact probabilities under the model distribution in the NRGBoost update. In practice, we would only be able to approximately estimate these with MCMC. We leave further analysis for future work.
A.3 NRGBoost Algorithm
In Algorithm 1 we provide a high-level overview of the training loop for NRGBoost.
Algorithm 1 NRGBoost training
Empirical training distribution $\hat{p}$, initial model distribution $q_{0}$, size of the sample pool $M$, fraction of samples to independently drop at each round $p_{refresh}\in[0,1]$, number of boosting rounds $T$, maximum number of leaves in each tree $J$, maximum probability ratio in each leaf $R>1$, shrinkage parameter $\gamma\in\left]0,1\right]$.
Energy function $f$.
$f\leftarrow\log q_{0}$
for $t\leftarrow 1$ to $T$ do
     if $t$ = 1 then $\triangleright$ No sampling needed since $q_{0}$ is normalized
         $\hat{q}\leftarrow q_{0}$
     else if $t$ = 2 then $\triangleright$ Initialize sample pool
         $\hat{q}\leftarrow\{\mathbf{x}_{i}\}_{i\in[M]}\sim_{iid}\exp f$triangleright By rejection sampling using $q_{0}$ as a proposal distribution
     else
         $\hat{q}\leftarrow$ Drop samples from $\hat{q}$ randomly with probability $p_{refresh}$
         $\hat{q}\leftarrow$ Keep each sample, $\mathbf{x}_{i}$, from $\hat{q}$ with probability given by Equation 12 using $\delta f_{t-1}$
         $\hat{q}\leftarrow$ Add $M-\lvert\hat{q}\rvert$ samples to $\hat{q}$ by Gibbs sampling from $\exp f$
     end if
     $\delta f_{t}\leftarrow\textrm{FitTree}(\hat{p},\hat{q};J,R)$triangleright Greedily grow tree using Equation 10 as a splitting criterion
     $\alpha_{t}\leftarrow\arg\max_{\alpha}\Delta L_{f}[\alpha\cdot\delta f_{t}]$triangleright Line search on training likelihood (Equation 43)
     $f\leftarrow f+\gamma\cdot\alpha_{t}\cdot\delta f_{t}$
end for
Appendix B Density Estimation Trees and Density Estimation Forests
Density Estimation Trees (DET), introduced by Ram & Gray (2011), model the density function as a piecewise constant function,
@@ Appendix B Density Estimation Trees and Density Estimation Forests
where the $X_{j}$ are given by a partitioning of the input space, $\mathcal{X}$, induced by a binary tree and the $v_{j}$ are the density values associated with each leaf that, for the time being, we will only require to be such that $q(\mathbf{x})$ sums to one. Note that it is possible to draw an exact sample from this type of model by randomly selecting a leaf, $j\in[1..J]$, given probabilities $v_{j}$, and then drawing a sample from a uniform distribution over $X_{j}$.
To fit a DET, Ram & Gray (2011) proposes optimizing the Integrated Squared Error (ISE) between the data generating distribution, $p(\mathbf{x})$ and the model:
@@
Noting that $\bigcup_{j=1}^{J}X_{j}=\mathcal{X}$, we can rewrite this as
@@
Since the first term in the objective does not depend on the model, this optimization problem can be further simplified as
@@
One can therefore grow a tree by greedily choosing to split a parent leaf, with support $X_{P}$, into two leaves, with supports $X_{L}$ and $X_{R}$, so as to maximize the following criterion:
@@
This leads to a similar splitting criterion to Equation 10, but replacing the previous model’s distribution with the volume measure $V$, which can be interpreted as the uniform distribution on $\mathcal{X}$ (up to a multiplicative constant).
@@
Often, generative models are trained to maximize the log-likelihood of the observed data,
@@
rather than the ISE. This was left for future work in Ram & Gray (2011) but, following a similar approach to the above, the optimization problem to solve is:
@@
This is, again, easy to solve for $v_{j}$ since it is separable over $j$ after removing the constraint using Lagrange multipliers. The optimal leaf values and objective are, in this case:
@@
The only change is, therefore, to the splitting criterion which should become:
@@
This choice of splitting criterion can be seen as analogous to the choice between Gini impurity and Shannon entropy in the computation of the information gain in decision trees.
Bagging and Feature Subsampling
Following the common approach in decision trees, Ram & Gray (2011) suggests the use of pruning for regularization of DET models. Practice has, however, evolved to prefer bagging as a form of regularization rather than relying on single decision trees. We employ the same principle for DETs, by fitting many trees on bootstrap samples of the data. We also adopt the common practice from Random Forests of randomly sampling a subset of features to consider when splitting any leaf node, in order to encourage independence between the different trees in the ensemble. The ensemble model, which we call Density Estimation Forests (DEF), is thus an additive mixture of DETs with uniform weights, therefore still allowing for normalized density computation and exact sampling.
@@ Appendix C Greedy Tree-Based Multiplicative Boosting
In multiplicative generative boosting, an unnormalized current density model, $\tilde{q}_{t-1}(\mathbf{x})$, is updated at each boosting round by multiplication with a new factor $\delta q_{t}^{\alpha_{t}}(\mathbf{x})$:
@@
For our proposed NRGBoost, this factor is chosen in order to maximize a local quadratic approximation of the log-likelihood around $q_{t-1}$, as a functional of the log-density (see Section 3). The motivation behind the greedy approach of Tu (2007) and Grover & Ermon (2018) is to instead make the update factor $\delta q_{t}(\mathbf{x})$ proportional to the likelihood ratio $r_{t}(\mathbf{x})\coloneqq\nicefrac{{p(\mathbf{x})}}{{q_{t-1}(\mathbf{x})}}$ directly. Under ideal conditions, this would mean that the method converges immediately when choosing a step size $\alpha_{t}=1$. In a more realistic setting, however, this method has been shown to converge under conditions on the performance of the individual $\delta q_{t}$ as discriminators between real and generated data (Tu, 2007; Grover & Ermon, 2018; Cranko & Nock, 2019).
In principle, this desired $r_{t}(\mathbf{x})$ could be derived from any binary classifier that is trained to predict a probability of a datapoint being generated (e.g., by training it to minimize a strictly proper loss). Grover & Ermon (2018), however, proposes relying on the following variational bound of an $f$-divergence to derive an estimator for this ratio:
$D_{f}(P\|Q_{t-1})\geq\sup_{u\in\mathcal{U}_{t}}\left[\mathbb{E}_{\mathbf{x}\sim p}\;u(\mathbf{x})-\mathbb{E}_{\mathbf{x}\sim q_{t-1}}f^{*}(u(\mathbf{x}))\right]\,,$ (60)
where $f^{*}$ denotes the convex conjugate of $f$. This bound is tight, with the optimum being achieved for $u_{t}^{*}(\mathbf{x})=f^{\prime}(\nicefrac{{p(\mathbf{x})}}{{q_{t-1}(\mathbf{x})}})$ if $\mathcal{U}_{t}$ contains this function. $\left(f^{\prime}\right)^{-1}(u^{*}_{t}(\mathbf{x}))$ can thus be interpreted as an approximation of the desired $r_{t}(\mathbf{x})$.
Adapting this method to use trees as weak learners can be accomplished by considering $\mathcal{U}_{t}$ in Equation 60 to be defined by tree functions $u=\nicefrac{{1}}{{J}}\sum_{j=1}^{J}w_{j}\mathbf{1}_{X_{j}}$, with leaf values $w_{j}$ and leaf supports $X_{j}$. At each boosting iteration, a new tree, $u^{*}_{t}$, can be grown to greedily optimize the lower bound in the r.h.s. of Equation 60, and setting $\delta q_{t}(\mathbf{x})=\left(f^{\prime}\right)^{-1}(u^{*}_{t}(\mathbf{x}))$. This means that $\delta q_{t}$ is also given by a tree with the same leaf supports and leaf values $v_{j}\coloneqq\left(f^{\prime}\right)^{-1}(w_{j})$, leading to the following separable optimization problem:
@@
Note that we drop the iteration indices from this point onward for brevity. Maximizing over $w_{j}$ with the $X_{j}$ fixed, we have that $w_{j}^{*}=f^{\prime}\left(\nicefrac{{P(X_{j})}}{{Q(X_{j})}}\right)$, which yields the optimal value
$J^{*}(X_{1},\ldots,X_{j})=\sum_{j}\left[P(X_{j})f^{\prime}\left(\frac{P(X_{j})}{Q(X_{j})}\right)-Q(X_{j})(f^{*}\circ f^{\prime})\left(\frac{P(X_{j})}{Q(X_{j})}\right)\right]\,.$ (62)
This, in turn, determines the splitting criterion as a function of the choice of $f$. Finally, the optimal density values for the leaves are given by
@@
It is interesting to note two particular choices of $f$-divergences. For the KL divergence, $f(t)=t\log t$ and $f^{\prime}(t)=1+\log t=\left(f^{*}\right)^{-1}(t)$, which leads to
@@
as the splitting criterion. The Pearson $\chi^{2}$ divergence, with $f(t)=(t-1)^{2}$, leads to the same splitting criterion as NRGBoost. Note, however, that for NRGBoost the leaf values for the multiplicative update of the density are proportional to $\exp\left(\nicefrac{{P(X_{j})}}{{Q(X_{j})}}\right)$ instead of the ratio directly. Table 4 summarizes these results.
@@
Another interesting observation is that a DET model can be interpreted as a single round of greedy multiplicative boosting, starting from a uniform initial model. The choice of the ISE as the criterion to optimize the DET corresponds to the choice of Pearson’s $\chi^{2}$ divergence and log-likelihood to the choice of KL divergence.